### v1

# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=gray!40

Abstract
Despite being a powerful concept, distributed shared memory (DSM) has not been made practical due to the extensive synchronization needed between servers to implement memory coherence. This paper shows a practical DSM implementation based on the insight that the ownership model embedded in programming languages such as Rust automatically constrains the order of read and write, providing opportunities for significantly simplifying the coherence implementation if the ownership semantics can be exposed to and leveraged by the runtime. This paper discusses the design and implementation of DistR, a Rust-based DSM system that outperforms the two state-of-the-art DSM systems GAM and Grappa by up to 2.64$\times$ and 29.16$\times$ in throughput, and scales much better with the number of servers.
Part of the work was done when Haoran Ma and Shi Liu interned at Alibaba Group.
1 Introduction
The concept of distributed shared memory (DSM) received significant attention during the early years of distributed computing systems. This era witnessed a plethora of pioneering efforts, as exemplified by seminal works such as [50, 51, 17, 18, 80, 10, 56, 32, 16, 62, 63, 61, 37]. DSM offers the power of parallel computing using multiple processors and machines and, more crucially, streamlines the development of distributed applications with a unified, contiguous memory view.
The initial enthusiasm for DSM was tempered by significant performance bottlenecks, primarily due to the low network speeds prevalent during its nascent stages. Recent advances in hardware and networking technologies [7, 78, 30, 19, 34, 43, 54, 64, 74, 39, 3, 12, 47, 41, 52, 66, 23] have revitalized the DSM explorations. Several new DSM systems [14, 60, 81, 87, 46, 77] were proposed in recent years to take advantage of these enhanced networks. However, these systems are still far from achieving satisfactory performance, exhibiting poor scalability and substantial slowdown compared to their single-machine counterparts. This is mainly due to the intensive synchronization operations needed to ensure memory coherence across servers.
State of the art. The majority of existing DSM systems [6, 46, 14, 87] adopt an approach to achieve data consistency by adhering to the following invariant: for each data block to be accessed, the block is either located on a single node with potential read and write access, or it is replicated across multiple nodes with each having read access only. Prior to a server attempting to access a block, a DSM system checks the state of the block, invalidates copies of that block on all other servers, and then transmits the block to the requesting server. This synchronization process necessitates multiple network round trips. Even with RDMA, the incurred latency is still orders of magnitude higher compared to a single local access, significantly degrading overall performance. Effectively reducing the number of synchronizations is, therefore, crucial for minimizing DSM overhead and rendering it feasible for real-world deployment.
A practical strategy to minimize synchronization overhead involves implementing high-level protocols to guarantee exclusive access for each server. For instance, Apache Spark [90] utilizes an immutable data structure known as a resilient distributed dataset (RDD) for distributed access. However, RDD only facilitates coarse-grained distributed access, limiting each server to accessing a distinct partition of an RDD. While increasing access granularity enhances performance, it comes at the expense of reduced generality—Spark is tailored for bulk processing of batch data and is incapable of supporting distributed applications requiring object-level accesses, such as social networks where objects of various types and sizes (e.g., images, connections, etc.) are created and manipulated upon each user request.
Insights. Our main observation is that synchronization overheads in existing DSM systems are introduced primarily due to the use of a generic approach that overlooks semantic information from programs. For example, many real-world concurrent programs are engineered with a single-writer-multiple-reader (SWMR) discipline to ensure correctness during concurrent operations. Leveraging such information can potentially eliminate the need to check the state of remote data blocks before accessing them, leading to dramatically improved performance. A major challenge is, however, how to expose such semantics in a sensible way so that the DSM system can see and act upon it.
One approach to convey such semantics, as demonstrated by AIFM [73] and Midas [68], involves exposing APIs that developers can invoke to specify program regions accessible only by a single writer. However, this process is cumbersome and error-prone, demanding a profound understanding of potential executions and involving substantial program writing. Our key insight in this endeavor is that the SWMR programming paradigm aligns seamlessly with ownership types, which have already been integrated into programming languages like Rust [75]. Rust is widely employed in the system community for dependable and secure implementation of low-level systems code.
Rust’s ownership type inherently upholds SWMR properties in any compiled Rust program. The fundamental concept behind the ownership type is that each value is ensured to have a single unique variable as its owner throughout the execution. While multiple references to a value are allowed, only the owner and mutable references can modify the value. Moreover, only one of these references is permitted to be used for modifying the value at any given point.
When developing a DSM system on top of an ownership-based language like Rust, SWMR semantics are inherently embedded in any Rust program by design. Effortlessly extracting such information becomes possible with basic compiler support, sparing developers from the need for code rewriting. Utilizing the SWMR semantics from the program leads to a considerably simplified process for accessing data in DSM. In the case of a write access, the ownership type ensures exclusive access to the data. Consequently, DistR can move the data to the requesting machine, performing the write there without explicitly invalidating its copies on other machines. In the case of a read access, data can be efficiently replicated to (and cached in) each requesting machine, benefiting from the compiler-provided assurance of freedom from concurrent writes.
This paper presents DistR, an efficient Rust-based DSM implementation that enables object-level concurrent accesses by leveraging the SWMR semantics made explicit by Rust’s ownership type. DistR automatically turns a single-machine Rust program into a DSM-based distributed version without requiring code rewriting. While extracting the ownership semantics appears straightforward, leveraging it to implement a distributed coherence protocol correctly and efficiently presents two main challenges.
The first challenge is how to manage memory correctly and efficiently. Rust’s ownership type system is inherently designed for a single-machine environment, where the memory address of an object remains constant post-creation. This assumption is disrupted in a distributed environment, where objects may be migrated or duplicated on different machines. Such actions can lead to the risk of dangling pointers, potentially breaking memory coherence.
To tackle these issues, DistR builds a global heap spanning multiple servers based on the idea of partitioned global address space [21]. Each object in the heap has a unique global address in the address space, which can be used for accessing the object from any server. DistR re-implements Rust’s memory management constructs to allocate objects in the global heap. Given that a server can have cached objects (to accelerate reads), DistR carefully crafts an ownership-based cache coherence protocol upon the global heap abstraction to achieve both memory coherence and efficiency (§4.1.1).
In a nutshell, our coherence protocol leverages the ownership semantics to eliminate the need for explicit cache invalidation. It allows multiple readers to fetch a copy of the object from its host server and cache it, but disallows any change to the global address and the value of the object. When a write access occurs, it must first borrow the ownership, at which point DistR moves the object in the global heap to a new address on the server issuing the write. The address change of the object automatically invalidates cache copies that use the stale address and triggers the subsequent readers to update the cache by fetching the object from its latest address.
The second challenge is how to support transparency in programming. Rust’s standard libraries and programs were originally built for running on a single machine, and they cannot deal with distributed resources in a cluster. For example, a Rust program running on server A cannot spawn a thread on another server B, let alone synchronize threads between A and B. To enable a Rust program to run as is under DistR, we provide distributed threading utilities by restructuring critical elements of the Rust standard library, including threading, communication channels, and shared-state locks (§4.1.2). Our adapted libraries offer the same interfaces, making them compatible with single-machine Rust programs, but internally invoke our distributed scheduler, which determines where to run the thread and facilitates cross-server synchronization. We built them atop the ownership-based memory model, enabling the DistR runtime to safely pass references of objects between threads and automatically fetch the value from the global heap upon dereferencing.
With our programming abstractions, a Rust application can start on a single server and gradually spawn its threads to other servers. Under the hood, DistR employs a runtime to manage distributed physical compute and memory resources for the application. The runtime runs as a process on each node in the cluster, and they work cooperatively for cross-server memory allocation and thread scheduling. The runtime prioritizes the current server for object allocation and thread creation, but it will schedule the resource allocation request to another server under memory pressure (§4.2.1). To make cluster-wise decisions such as deciding the target server for global memory allocation and thread creation, DistR has a global controller that is launched together with the application. The global controller communicates with DistR runtime on each node to collect resource usage information and applies adaptive policies to achieve load balance (§4.2.2).
Results. We evaluated our system on four real-world applications in an eight-node cluster. Our evaluation demonstrated an average of 2.02$\times$ and 9.48$\times$ (up to 2.64$\times$ and 29.16$\times$) speedup compared with two state-of-the-art DSM systems GAM and Grappa, respectively. Furthermore, DistR incurred a mere 2.42% slowdown compared to the original Rust program on a single machine with sufficient resources.
2 Background in Ownership
Over the past decades, numerous programming languages have been designed to provide safe memory management and data sharing. At the core of such a design is often a tradeoff between memory abstraction level and management efficiency. The ownership concept, and the Rust programming language built upon, are considered promising solutions that achieve a sweet spot between abstraction and efficiency. This section provides an overview of these techniques and explains how ownership can benefit DSM implementations.
Ownership Type. The ownership model has a long history in pursuit of memory-safe language designs and type systems [8, 28, 44, 58, 83, 9]. It has also inspired many systems for safe and efficient resource management [42, 59, 13, 88]. At a high level, ownership enhances a language’s type system in a way that guarantees the memory and thread safety of a program with type checking done at compile time. The ownership model encompasses a range of concepts, among which the most important are lifetimes and borrowing.
An ownership-based type system uses lifetimes to control the allocation/deallocation of objects. It enforces that each object must have one and only one owner at a time. This allows the compiler to statically track an object’s lifetime via its owner, and immediately deallocate the object once its owner goes out of scope, preventing memory leaks without using garbage collection that can introduce disruptive pauses to program execution.
To access an object, a program can create a reference from its owner, but the reference must “borrow” the permission from the owner, and “return” it to the owner after the access. Specifically, the type system allows the creation of multiple immutable references to an object from its owner for concurrent reads but prohibits any write with these references. It allows only one mutable reference to the object only when no other (mutable or immutable) references exist. Through borrowing, the ownership type disallows simultaneous writers and hence prevents data races. In addition, references must return the borrowed permission when they go out of scope. For any program that demonstrates type soundness, the type checker guarantees that references to an object can only reside within the object’s lifetime; the object can be safely and automatically deallocated when its owner goes out of scope, by which time it has already lost all its references.
Finally, ownership can be transferred from one owner to another—e.g., at a function call, the creation of a thread, or message passing (i.e., via channel). However, the type system enforces that ownership transfer must occur in the absence of “borrowing”. In other words, no other references can exist in scope when transferring the ownership, preventing data races during ownership transfers.
The guarantees provided by the ownership model with respect to object lifetime and data sharing can be summarized with the following four invariants:
  1. 1.
    Singular Owner: each value has one single owner at any time (which must also belong to one single thread).
  2. 2.
    Safe Borrowing: All references are created from the owner; permission borrowing and returning guarantees that references that can be used to access the object must be valid.
  3. 3.
    Single Writer: Each object allows one mutable reference at most, and it cannot coexist with any other references in the same scope.
  4. 4.
    Multiple Reader: Multiple references are permitted only when all of them are immutable.
The last two invariants are commonly called the single-writer-multiple-reader (SWMR) property in the DSM literature [57].
Figure 1: State machine for Rust’s ownership-based memory model.
Rust Language. Rust offers a practical implementation of ownership and is designed with a range of zero-cost abstractions for efficient fine-grained resource management. Figure 1 depicts the state machine for Rust’s ownership-based memory model. At a high level, this model restricts that the owner is always in the O (owned) state, and transitions between M (modified), S (shared), and I (invalid) must go through the O state [^1] A transition from M to S is also possible as an optimization in Rust.. Clearly, a distributed implementation of this approach avoids broadcasts or snooping, and only requires peer-to-peer message passing.
[…]
2impl Accumulator {
3 pub fn add(&mut self, delta: &i32)->i32 {
4 *self.val += *delta;
5 *self.val
6 }
[…]
8fn main() {
9 // Allocates two integers in the heap.
10 let val: Box<i32> = Box::new(5); // val is an owner.
11 let mut b: Box<i32> = Box::new(0); // b is an owner.
12 // Ownership is transferred from val to a.val
[…]
14 { // Only one mutable reference is allowed.
15 let mutr: &mut i32 = &mut *b;
16 // No other reference is allowed now.
17 /* let another_r = &*b; */ // COMPILE ERROR!
18 *mutr = 10; // b == 10
[…]
20 { // Multiple immutable references are allowed.
21 let (b_r1, b_r2): (&i32, &i32) = (&*b, &*b);
22 // Mutable reference is prohibited now.
23 /* let b_mutr = &mut *b; */ // COMPILE ERROR!
24 // Passing by references won’t transfer ownership.
25 let sync_add = a.add(b_r1); // a.val == 15
26 let sync_add = a.add(b_r2); // a.val == 25
[…]
28 {// Ownership of a and b is moved to the new thread.
29 // No reference should or can borrow a or b now.
30 let async_add = thread::spawn(move ||
31 a.add(&*b) // a.val == 35
32 ).join(); // lifetime of a and b ends
33 // Current thread cannot access a and b anymore.
34 /* println!("{}", a.val); */ // COMPILE ERROR!
[…]
36}
Listing 1: A simple accumulator implementation in Rust.
Listing 1 exemplifies a simple accumulator implemented in Rust (Lines LABEL:lst:simple-acc:acc-sttLABEL:lst:simple-acc:acc-end). The Accumulator struct keeps an integer val and exposes an interface add to increment the value. Rust uses a smart pointer type Box<T> to store values on the heap; this pointer serves as the initial owner of the referenced value, as shown in Line 10 and 11. Line LABEL:lst:simple-acc:inst-acc instantiates Accumulator a, where the ownership is implicitly transferred from val to a.val during its initialization. Rust allows the creation of mutable and immutable references to access the value. For example, Lines 14LABEL:lst:simple-acc:mut-end create a singular mutable reference (&mut) to b and set its value to 10. Similarly, Lines 20LABEL:lst:simple-acc:immut-end create two immutable references (&) to b and add them to a via two function calls. Note that passing references as arguments in function calls does not transfer their ownership.
Finally, Rust allows spawning new threads for concurrent programming, as shown in Lines 28LABEL:lst:simple-acc:conc-end. A new thread is created via thread::spawn, where the use of move captures a and b in the current scope and transfers their ownership to the newly spawned thread. Rust performs shallow copying for inter-thread communication, where only the pointers stored in a and b are transferred to the child thread while the actual values on the heap are not moved. Rust guarantees memory safety of a and b by tracking their ownership. At Line 32, when the child thread finishes its closure (i.e., not necessarily after join), and a and b exit the scope (to which their ownership belongs), their lifetimes terminate and Rust deallocates them from the heap.
3 Motivation
DSM was proposed to eliminate the barrier of distributed programming by offering the same memory consistency model as single-machine shared memory. The core of its design is a software-based cache coherence protocol, which mimics a hardware-based approach on multi-core CPUs and synchronizes memory states on different servers by sending control messages between them. However, it is notoriously hard to implement cache coherence efficiently at the software level due to the high communication latency between physically disjointed servers.
High Synchronization Overheads for Coherence. To gain a high-level understanding of how much improvement can be achieved by improving the cache coherence protocol, we performed an analysis by running a real-world application DataFrame [67] with a state-of-the-art DSM system GAM [14] with a fast network. We first ran Dataframe on a single server with 16 CPU cores and 64GB memory. We then ran it with GAM on eight servers connected by a 40Gbps Infiniband network by evenly distributing the same amount of resources to eight servers (i.e., each server uses 2 CPU cores and 8GB memory). Our experiments show a 2.4$\times$ slowdown when Dataframe runs on eight servers.
[…]
4 Design
Figure 2: Design overview of DistR.
DistR is an efficient DSM framework atop the Rust programming language. As shown in Figure 2, it consists of Rust-based programming abstractions for DSM (§4.1) and a runtime (§4.2) that manages distributed physical resources.
DistR is compatible with standard Rust. Listing 2 illustrates how the accumulator (shown in Listing 1) runs on DistR distributively without requiring code rewriting. The program starts running on a single machine A and the DistR runtime gradually allocates its memory and spawns new threads on different machines. Specifically, Lines 10LABEL:lst:dist-acc:line:acc-end create Accumulator a and b where a.val and b are in the global heap. We use a global allocator to allocate objects in the global address space and hence these objects may be allocated on a different server. Line 15 synchronously adds b to a by fetching both values a.val and b to A’s local memory (if they are allocated somewhere else). Line LABEL:lst:dist-acc:line:remote-add spawns a new thread and ships the function closure to perform add asynchronously. This thread will be scheduled on a different server B if A’s compute power has been saturated. In this case, DistR performs shallow copying and only ships the pointers stored in a and b to B without actually moving objects in the global heap. The newly-created thread relies on the DistR runtime to detect data locations and fetch objects upon dereferencing.
1// Unmodified Rust code.
2pub struct Accumulator { pub val: Box<i32>, }
3impl Accumulator {
4 pub fn add(&mut self, delta: &i32)->i32 {
5 *self.val += *delta;
6 *self.val
7 }
8}
9fn main() {
10 // Allocates two integers in the distributed heap.
11 let val: Box<i32> = Box::new(5);
12 let b: Box<i32> = Box::new(10);
[…]
14 // a.val and b will be fetched to local.
15 let local_add = a.add(&*b); // a.val == 15
16 // Only refs to a and b are shipped to remote.
[…]
18 a.add(&*b)).join(); // a.val == 25
19}
Listing 2: DistR seamlessly transforms an unmodified accumulator implemented in Rust into a distributed version.
4.1 DistR Programming Abstraction
DistR provides each thread with a local stack and abstracts distributed memory as a shared global heap. Each server allocates thread stacks and backs one partition of the global heap with its physical memory. DistR re-implemented core memory management constructs including Box, &, and &mut for transparent heap access. This approach hides the complex details of memory allocation/deallocation, moving objects, and coherence maintenance (§4.1.1). DistR supports distributed threading and synchronization by adapting Rust’s standard libraries atop the core language constructs (§4.1.2). Furthermore, DistR offers affinity annotations that allow developers to build more efficient applications by expressing data affinity semantics (§4.1.3).
4.1.1 Memory Management
Figure 3: The address space layout of DistR. The stack is private to each thread but they share an aligned address space to ease migration, while the heap is globally shared and partitioned across servers.
Next, we discuss how DistR (re)implements the memory-related language constructs in Rust to achieve memory safety and memory coherence.
Address Space. As shown in Figure 3, DistR maintains an identical address space layout on all servers. It exposes distributed memory as a coherent shared heap to applications. Embracing the idea of partitioned global address space (PGAS) [21], it partitions the heap space and assigns each server a unique address range. The stack, in contrast, is private to each thread. However, DistR aligns the stack space on each server and pads stacks to avoid overlapping. This streamlines thread migration between servers as it allows a thread to keep its private stack address unchanged when being moved.
Coherence Protocol in a Nutshell. For efficiency, DistR employs a call-by-reference model for newly created threads. Upon creation of a thread, the DistR runtime only passes references or Box pointers to objects to the newly created thread. Upon dereferencing, objects are fetched to the server where the thread is executed.
When a read access of an object is issued on a server, our runtime simply fetches a copy of the object from its hosting server and places it in its local cache. As a result, multiple copies of the same object may exist on different servers. This allows multiple servers to read the object at the same time from their respective cached copies. Fetching a copy of the object for read does not change the object’s address in the global space. When a write access occurs on an object, the server issuing the write must first obtain the object’s write access permission through a mutable borrow. Our reimplementation of mutable borrow (discussed shortly) moves [^2] The term “copy” is used to describe the process of adding an object into the cache without changing its global address. The term “move” means relocating the object into a server’s heap partition, which requires changing its global address. the object in the global heap to a new address that belongs to that server. In doing so, the object’s cached copies on other servers are automatically invalidated without sending explicit invalidation messages—subsequent reads on these servers must obtain an immutable reference to the object through an immutable borrow from its owner pointer, which has been updated to the new address immediately after the mutable borrow returns. Upon identifying the owner’s address change, each immutable borrow would direct a server to fetch a fresh version of the object from the new address as opposed to relying on a stale copy residing in its cache.
Note that this is a general protocol that covers the case that the object is on the same server that issues the write—as long as the server moves the object into a different location in the global heap, no other servers can read the stale copies of the object. However, this is not efficient as each local write requires moving the object to a new address. To address this inefficiency, DistR employs a pointer-coloring technique, inspired by the designs of many concurrent garbage collectors [1, 53]. Discussed at the end of this subsection, this technique offers a more efficient solution for handling local writes.
Figure 4: DistR repurposes Rust pointers and references to contain a global heap address and an extension field for its coherence protocol.
Pointer Layout. In order to support this protocol, each pointer must remember not only the object’s global address, but also the address of the cached copy in a server’s local cache (to avoid redundant remote fetches). As such, we modify Rust’s pointer structure, as illustrated in Figure 4. DistR internally extends each Rust Box pointer and reference with an additional 64-bit field, which is used differently for read and write access. At a high level, the field records the address of the cached copy for faster read accesses; for write accesses, this field records the address of the object’s owner for post-write synchronization. Additionally, DistR reserves the highest 16 bits in the global address field as “color” bits. These bits record the version number of the pointer and play a crucial role in DistR’s efficient handling of local writes.
Next, we discuss how DistR reimplements Rust’s ownership operations to realize the distributed coherence protocol. For ease of presentation, this subsection focuses on a simplified version of the protocol. A complete coherence protocol and its proof of memory coherence are available in Appendix B.
Input: A mutable reference $m$ containing a global address $m.g$ and the owner address $m.o$.
Output: A local memory address to be written to.
1
2 Function DerefMut($m$):
3       if $\neg$¬IsLocal($m.g$) then
4             $m.g\leftarrow$ Move(ClearColor($m.g$))
5
6      return ClearColor($m.g$)
7
8
9Function DropMutRef($m$):
10       $c^{\prime}\leftarrow$ GetColor($m.g$) $+1$
11       Write($m.o$, AppendColor($m.g$, $c^{\prime}$))
12
Algorithm 1 Access logic for mutable references.
Mutable Borrow. Mutable borrow creates a mutable reference that holds exclusive access to the referenced object for writing. Algorithm 1 outlines the procedures for both dereferencing and dropping a mutable reference. When performing dereferencing, DistR first checks the object’s location (Line 1) and performs direct access if the object’s address belongs to the heap partition of the machine A that executes the access. Otherwise, DistR moves it to A’s heap partition (as opposed to caching it) (Line 1). The move, conducted in the following three steps, changes the object’s global address. DistR (1) copies the object into A’s heap at an address $p$, (2) updates the mutable reference with the address $p$, and (3) asynchronously requests the remote server that previously stored the object to deallocate the original object.
A challenge arises with its original owner Box, which now becomes a dangling pointer, pointing to an invalid memory location. Fortunately, the integrity of the system is maintained by the single-writer invariant (referenced as Invariant 3). This invariant ensures that while the mutable reference remains alive, no other entity, including the original owner, can access the data. To ensure correctness, when this new reference is dropped, DistR synchronously updates the original owner Box, redirecting it to the new address $p$ (Line 1). As a result, the original owner always possesses the latest view of the object. Additionally, all modifications made through this mutable reference are visible in all subsequent accesses, as they necessitate borrowing permission from the updated owner Box. The single-writer invariant also eliminates the possibility of simultaneous updates to the owner, ensuring that updating the owner is free from concurrency issues.
Input: A shared immutable reference $r$ containing a global address $r.g$ and a local copy address $r.l$, and a local cache hashmap $\mathit{H}$.
Output: A local memory address for reading.
1
2 Function Deref($r,H$):
3       if IsLocal($r.g$) then
4             return ClearColor($r.g$)
5
6      else
[…]
8                   Atomic {
9                   if $r.g\in H$ then
10                         $\langle l^{\prime},cnt\rangle\leftarrow$ GetEntry($H$, $r.g$)
11                         $r.l\leftarrow l^{\prime}$
12                         UpdateEntry($H$, $r.g$, $\langle l^{\prime}$, $cnt+1\rangle$)
13
14                  else
15                         $r.l\leftarrow$ Copy(ClearColor($r.g$))
16                         InsertEntry($H$, $r.g$, $\langle r.l,1\rangle$)
17
18                  }
19
20            return $r.l$
21
22
23
24Function DropRef($r,H$):
25       if $r.l\neq\text{Null}$ then
26             Atomic {
27             $\langle l^{\prime}$, $cnt\rangle\leftarrow$ GetEntry($H$, $r.g$)
28             UpdateEntry($H$, $r.g$, $\langle l^{\prime}$, $cnt-1\rangle$)
29             }
30
31
Algorithm 2 Access logic for immutable reference.
Immutable Borrow. Immutable borrowing allows concurrent reads to the same object from immutable references on the same or different servers. As detailed in Algorithm 2, DistR handles the dereferencing of immutable references by first checking the object’s location (Line 2). For remote objects, DistR creates a local copy in the per-node read-only “cache” and records its local address in the reference’s extension field (see Figure 4). This preserves the original global address of the object, ensuring that any new immutable reference—whether it is derived from the owner Box or from another immutable reference—can always access the original object from the global heap.
As opposed to being a separate memory space, our “cache” provides a “virtual” aggregation of all local copies maintained on each server. These copies reside in the regular heap, managed by a per-node hashmap $H$. This hashmap maps each global address to a pair of its local address and the number of local immutable references to the local copy. To prevent redundant copies of an object on the same server, DistR checks the hashmap $H$ before creating a new local copy (Line 2). If a local copy is already present, DistR increments its reference count in $H$ and updates the extension field in the immutable reference to point to this copy (Lines 22). If no existing copy is found, a new one is created (Lines 22). Since the hashmap uses objects’ global addresses as keys, if an object has been modified by another server since its last read, its global address must have changed, making cache lookup fail even if a (stale) local copy exists.
DistR actively updates the reference count of each local copy when an immutable reference is either dereferenced or dropped, as outlined in Lines 2 and 2. Utilizing these counts, the DistR runtime periodically scans the “cache” and lazily reclaims unreferenced copies (i.e., those with a zero reference count) under memory pressure (§4.2.1). This mechanism, in conjunction with the safe borrowing invariant (2), prevents the local cache from memory leaks or illegal accesses.
Owner Access without Borrow. DistR treats a direct memory access via the owner Box as a pair of mutable/immutable borrow and return. Depending on the reference type, DistR uses the extension field of Box accordingly and executes the read/write dereferencing logic. A special case arises when a mutable owner is immutably borrowed and becomes immutable until all borrowed references return. In this case, the owner can only cache the object during the borrow and delay the move until the borrow finishes. This would not create any correctness issues because the owner cannot be used for write access during this period.
Ownership Transfer. Similar to Rust, DistR does not move the actual value during the transfer and only copies the Box pointer. DistR additionally checks and resets the pointer’s extension field and frees the cached copy in the executing machine’s cache to avoid cache leakage.
Memory Deallocation. Like Rust, DistR tracks the lifetime of an object via its owner. Given that ownership transfer is implemented by only evicting the cached copy of the object (without changing its global presence), the memory safety of DistR’s global heap is preserved by the singular owner invariant (1). In other words, DistR still guarantees that when an object’s owner goes out of scope, the object must be unreachable (and dead) and can be safely deallocated.
Consistency Model. Our protocol, together with Rust’s ownership model, offers sequential consistency for cross-server memory accesses in safe Rust programs (i.e., following the original Rust, no guarantees can be provided when Rust Unsafe is used), which is a strong consistency order. Therefore, it allows any safe Rust program to preserve its memory consistency on DSM. Sequential consistency necessitates a coherent memory system, requiring not only the SWMR invariant but also the data-value invariant [57]. In simple terms, the data-value invariant requires that the latest write to a value is immediately visible to subsequent readers. As discussed earlier, DistR’s protocol moves an object upon a write and updates the owner immediately. Therefore, the latest value is globally visible after each mutable borrow finishes. Subsequent read accesses, either in the Owned state or the Shared state, are hence guaranteed to see the moved object and read its latest value.
1 Function GetColor($g$):
2       return $g\gg 48$
3
4
5Function ClearColor($g$):
6       return $g$ $((1\ll 48)-1)$
7
8
9Function AppendColor($g$, $c$):
10       return ClearColor($g$) | ($c\ll 48$)
11
Algorithm 3 Utility functions for pointer coloring.
Optimizing for Local Writes. A special case is that a server issues a write to an object that resides in its own heap partition. While the coherence protocol still guarantees safety, requiring moving an object in its local heap each time it is written clearly brings inefficiencies. To optimize for local writes, DistR adopts a pointer-coloring method, inspired by the design of concurrent garbage collectors in a managed runtime system such as JVM [1, 53]. Several utility pointer coloring functions are shown in Algorithm 3 which are used when dereferencing and dropping a reference. We reserve the first 16 bits of a global address as a “color”. The color value stored in the object’s owner gets incremented upon the expiration of a mutable reference, as detailed in Lines 11 in Algorithm 1. Any subsequent immutable borrow would look up the cache with the object’s global address. Even if the actual address remains the same, its color changes if a write has occurred. As such, the lookup would not return any stale copy from the local cache.
The 16-bit color field may overflow when the pointer keeps being borrowed for local writes on the same server. DistR implements a move-on-overflow strategy that moves the object to a new address and resets its color to zero once the maximum color value is reached ($2^{16}$), thereby preventing overflow and maintaining system integrity and performance.
Writing Unsafe Code in DistR. Rust allows developers to bypass compiler safety checks and write unsafe code for low-level operations such as accessing raw pointers and mutating shared variables at their own risk [45, 70]. Since DistR relies on SWMR semantics enforced by Rust’s ownership types, DistR ensures consistency and memory safety only in the “safe” Rust code. DistR does not cache objects in unsafe code but allows developers to implement their own cache. Developers must ensure that they do not violate consistency in unsafe code blocks where type safety is not enforced. This caution mirrors practices in other managed languages, like native code in Java and unsafe code in C#. To assist developers, DistR offers primitives such as dalloc, dread, and dwrite for managing data on the global DSM heap.
4.1.2 Adapting Rust Standard Libraries
To further reduce the barrier for programs to run distributively, we reimplement several standard Rust libraries atop DistR’s core memory constructs covering four categories: threading for distributed computation (std::thread), inter-thread channel for communication (std::sync::mpsc), reference-counted pointers for ownership sharing (std::sync::Rc and std::sync::Arc), and shared-state locks for concurrency control (std::sync::Mutex and std::sync::atomic).
Threading. DistR’s threading library enables Rust threads to run distributively with two major adaptations. First, it enables distributed thread launching by re-implementing the spawn interface. Internally, it captures the thread body as a closure during compile time and forwards it to the runtime. During execution, the runtime launches the thread according to each server’s load (details in §4.2.1). Second, DistR performs implicit ownership transfers between the parent and the child threads at the start or the end of the child thread execution. Thanks to the distributed ownership transfer support provided by DistR’s memory model, the implementation in the threading library is hidden from developers and preserves type soundness and memory safety. Additionally, DistR is compatible with advanced thread utilities such as thread::scope, which allows for the spawning of scoped threads that can borrow non-static data. These utilities ensure that all threads are joined at the end of their scope and can internally utilize DistR’s functions for spawning and joining threads, thus extending their applicability to the distributed setting.
Inter-Thread Channel. DistR extends Rust’s channel to connect two distributed threads for message passing. DistR internally builds a network-based message queue for cross-server messages. Benefiting from the shared global heap, Box pointers and references can be safely copied and remain valid across servers. Therefore, the sender can push an object into the channel as is without serialization, even if it may contain Box pointers. DistR forwards the object binary bytes to the receiver over the network, and the receiver can recover the object from the binary by direct type conversion without deserialization.
Ownership Sharing. Rust allows multiple owners to share an object via reference-counted smart pointers, which count the number of live owners. In this case, smart pointers only have read access, and the object lifetime terminates when all owners die and the reference count hits zero. DistR does not require special treatment for Rc as it only allows ownership sharing inside a single thread. For Arc which shares ownership among multiple threads, DistR handles it in a similar way to immutable references with on-demand local caching and lazy eviction.
Shared-State Concurrency. Rust supports shared-state concurrency, primarily through its atomics and mutexes, where threads commonly share an atomic-typed value or one mutex via ownership sharing (i.e., Arc). Unfortunately, the ownership model cannot type check concurrent read/write to shared states. Hence, Rust relies on an unsafe implementation in its standard library. §4.1.1 already provides a general discussion on writing unsafe code in DistR, and here we focus on DistR’s implementation for distributed shared states.
Shared states create a unique challenge for DistR, as they may be replicated on multiple servers and those states must be synchronized among these servers. For example, an Arc<AtomicBool> may be replicated across different servers and used independently, causing multi-version issues if not synchronized properly. DistR addresses this inconsistency by allocating the actual value on the global heap and storing only the Box pointer in atomic types. This design allows atomics to be freely moved or replicated across servers while keeping a single version of the actual value. To synchronize concurrent operations on atomics, DistR adapts methods of atomic types to forward the operation as a message to the server storing the actual value, which serializes all operations with unsafe logic similar to the original Rust to guarantee atomicity and memory consistency. Similarly, DistR implements Mutex by allocating its metadata and owned object on the global heap and leaving only Box pointers in the mutex struct. Concurrent operations on mutexes are serialized on the server storing the mutex.
4.1.3 Affinity Annotations
To further improve performance, DistR allows developers to provide optional data affinity semantics via annotations. These annotations are useful for many datacenter applications that make extensive use of object-oriented data structures that require pointer-chasing to access. For instance, Memcached [55] uses a chained hash table where each hash bucket stores its KV pairs with a linked list. To find one KV pair from a bucket, Memcached has to iterate the linked list following the node pointers. However, frequent pointer chasing is unfavorable in a distributed setting, where each pointer dereference incurs additional runtime checks and potential cross-server traffic. It would be beneficial for the runtime to colocate them on the same server and schedule the computation there.
[…]
3impl List {
[…]
5 let mut total: i32 = 0;
6 if let Some(r) = &self.head {
7 let mut node = &**r; // Fetch whole list to local.
8 loop { // Iterate every list node.
9 // Accessing node is guaranteed local.
10 total += (*node).val;
11 if let Some(next) = &node.next {
12 node = &**next;
13 } else { break; }
[…]
15 }
16 total
[…]
18}
Listing 3: A linked list implementation with TBox in DistR. The use of TBox ties list nodes one by one. Iterating a list will fetch all nodes together (if they are on another server), and henceforth accessing any node is guaranteed local.
Data-Affinity Pointer. To expose data affinity for clustered placement, DistR includes a new pointer type TBox for developers to “tie” a heap object to its owner. TBox shares the same interfaces as the ordinary Box and can be used as a drop-in replacement for Box. However, TBox enforces that the pointed-to object must reside on the same server as its owner. In other words, when its owner object is copied or moved, the object referenced by TBox will be copied or moved as well. TBox can be used in a nested manner to allow a large union of objects to be co-located. The DistR runtime fetches (i.e., copies or moves) them together in a single batch, leading to fewer network round-trips and higher throughput. TBox can also be assigned to and owned by a stack variable, in which case the referenced object is pinned onto the heap partition of the server that hosts the stack and cannot be moved. Dereferencing a TBox is thus guaranteed to be a local access—DistR optimizes it by skipping the runtime check.
Listing 3 presents a linked list implementation using TBox. Our linked list uses TBox (Line LABEL:lst:list:tbox-dec1) to specify the data affinity between consecutive nodes. As a result, all list nodes are chained with TBox, forming an affinity group. Line LABEL:lst:list:sum-sttLABEL:lst:list:sum-end define a sum function that calculates the total sum of all node values. Assuming the list is non-empty, Line 7 dereferences the pointer to the head node, and the DistR runtime checks the location of the head node and fetches the entire list of nodes together if they are not local. Next, accessing each node inside the loop body (Line 8LABEL:lst:list:loop-end) is guaranteed local and hence skips runtime checks. Compared to using Box directly, TBox makes both data fetching and accessing more efficient.
1fn main() {
2 let val: Box<i32> = Box::new(5);
3 let mut a = Accumulator{val};
4 let remote_add = spawn_to(a.val, move ||
5 a.add(10)).join(); // a.val == 15
6}
Listing 4: A distributed accumulator can leverage spawn_to to offload a thread to the server where a.val locates.
Data-Affinity Thread. To expose the affinity between data and computation for thread scheduling, DistR extends its threading library with a spawn_to interface. spawn_to mirrors the ordinary spawn interface to spawn a new thread but takes an additional Box pointer argument, which indicates where the thread should be created. The runtime checks where the Box points to and creates the new thread on that same server. A common practice to use spawn_to is to pass the mostly-accessed object as the location indicator. Listing 4 presents how the distributed accumulator (shown in Listing 2) can use spawn_to to offload a thread to the same server as a.val resides. Line 5 hence performs local dereference to a.val inside a.add().
4.2 DistR Runtime System
DistR’s runtime system is the core component that manages memory and compute resources. It includes a runtime library (§4.2.1) that is linked to each application and launched on each server and a cluster-wise global controller (§4.2.2).
4.2.1 Application-Integrated Runtime
The runtime library consists of a communication layer to support inter-server coordination and data transfer, a heap allocator to manage the heap partition and the read-only cache, and a thread scheduler to launch and schedule application threads.
Communication Layer. The DistR runtime uses its communication layer to support the cache coherence protocol and cross-server memory accesses. The communication layer consists of a control plane and a data plane, both built with RDMA. The control plane leverages two-sided verbs to send and receive small control messages, and the receiver can perform the coherence logic upon receiving the message to minimize the end-to-end latency. The data plane, in contrast, is specialized for bulky data transfer with one-sided verbs. It fetches an object as a whole with a single RDMA message upon pointer dereferencing without interrupting the target server, minimizing both latency and CPU usage.
Heap Allocator. The DistR runtime provides standard memory allocation interfaces and always returns global addresses to the upper-level language abstractions. It prioritizes local memory allocation as long as the local heap partition has sufficient space. This strategy improves data locality by colocating an object with its creating thread.
When the local heap partition is depleted, DistR queries the global controller and allocates memory on the most vacant server. For remote memory allocation, it forwards the request to the target server by sending a message through the communication layer and returns the allocated address to the user. Memory deallocation follows a similar logic but it bypasses the controller and finds the server directly via the object’s global address. The allocator does not reserve separate space for the local cache. Instead, it manages the cache as part of the local heap partition and always allocates cached entries in the local heap partition. Under memory pressure, the allocator will scan the local cache and evict entries that are no longer used (i.e., reference count hits zero).
Thread Scheduler. The DistR thread scheduler runs in the user space and schedules threads locally to maximize CPU utilization. It also provides thread migration functionalities, facilitating the global controller to balance load between busy and vacant servers.
The scheduler represents a newly created user thread as a closure, which includes a function pointer and a set of initial arguments (i.e., references). It collaborates with the global controller to allocate a unique stack space for a thread (see Figure 3), and starts the thread by executing the closure.
The scheduler adopts the method of cooperative multitasking and context switches between threads non-preemptively. A running thread yields its control flow proactively when developers call await or reactively upon long-latency operations. Similarly to other systems [60, 65, 82], our scheduler handles context switches as function calls, which allow DistR to save only a few registers per thread.
The scheduler supports creating/migrating a thread to another server as well. To migrate a thread, DistR sends its function pointer, the saved register state, and its stack to the target server. Because each thread reserves its stack address range globally, DistR can copy the stack across servers without changing its address. Therefore, the thread can be easily resumed by directly calling the function pointer with the saved register state on the target server. DistR generates code for state transmission during the compile time for the scheduler to call upon thread migration.
4.2.2 Global Controller
The controller runs as a daemon process on the machine where the program is launched. It manages cluster resources and coordinates memory allocation and thread migration. It periodically pings each server to probe and record its resource usage (CPU and memory). It controls resource allocation in cooperation with the DistR runtime on each server. When allocating memory or creating a thread, the runtime will first query the controller, which chooses a target server following adaptive policies (discussed later), and then coordinate with the runtime on the target server to perform the actual operation. The controller also maintains a global table to track the location of each thread; the table is queried and updated by the scheduler when migrating a thread.
During program execution, servers may run into imbalanced loads when objects get relocated or new threads are created. DistR balances the load of each server by migrating threads from the busy server to less occupied ones, following an adaptive policy to minimize cross-server memory accesses. If a server is about to run out of memory (>90% memory usage), the controller keeps migrating the thread that consumes the most local heap memory until the pressure is resolved. If the server is under compute congestion (>90% CPU usage), the controller migrates threads that frequently access remote objects. The thread is then moved to the server it accesses the most unless the target server is also overloaded, in which case it moves to a vacant server instead.
4.2.3 Fault Tolerance
In DistR, the global heap can be replicated to tolerate failures. Replication creates copies for each heap partition at the same virtual address on backup servers. Threads, in contrast, are not replicated for efficiency and are only executed with the primary global heap. To maintain a synchronized view between the primary heap partition and its backup copy, a thread must additionally write back to the backup partition after each mutable borrow. However, our insight is that the thread can batch modifications to an object and delay the write-back until the object ownership is transferred to another server, which is the time point that the object becomes visible to threads on other servers. When a server with a primary heap partition fails, the controller will automatically promote its backup server to the primary and add a new backup server.
5 Implementation
The majority of DistR was implemented in Rust except for its communication library which is in C. We implemented DistR’s core language constructs as three Rust types (i.e., struct): Ref<T>, MutRef<T>, and DBox<T>. They serve as the counterpart for the original Rust &T, &mut T, and Box<T>, respectively. We implemented the coherence protocol with traits on these types, including Copy, Clone, and Drop, which are automatically embedded into the program source code and executed when references/pointers are created or destroyed. To support unmodified Rust programs, we changed the Rust compiler and added additional compilation passes to transform Rust references and Box pointers into corresponding types in DistR.
Our communication layer links libibverbs directly for fast and kernel-bypassing RDMA networking. We implemented a low-level C library that covers basic connection establishment and exposes high-level Rust interfaces for various RDMA verbs, including RDMA_READ, RDMA_WRITE, RDMA_SEND, RDMA_RECV, ATOMIC_FETCH_AND_ADD, and ATOMIC_CMP_AND_SWP. We primarily utilize one-sided READ and WRITE verbs for data transfers between servers, as they outperform the two-sided SEND/RECV counterparts—one-sided operations bypass the CPU and OS at the receiver side, whereas two-sided operations require the receiver to pre-post RECV verbs and await notification upon message arrival. For instance, when a remote object is accessed via mutable references, DistR copies the object to local memory using the READ verb. Upon dropping the reference, DistR updates the original owner Box to reflect the new address, a process executed using the WRITE verb. Conversely, two-sided SEND/RECV verbs are utilized for control message exchanges, such as establishing connections across servers. Atomic verbs ATOMIC_FETCH_AND_ADD, and ATOMIC_CMP_AND_SWP are primarily utilized for implementing shared states (e.g., atomic types and mutexes). DistR uses the RC (reliable connection) transport type to ensure reliable transmission and strict message ordering.
Our heap allocator implementation piggybacks Rust’s original allocator and aligns its virtual address range with the heap partition range. Our thread scheduler was built upon Tokio [82] for its efficient user thread and cooperative scheduling integration. The global controller is responsible for managing all threads in the cluster and padding their stacks to avoid address overlapping.
6 Limitations
DistR’s design has three limitations. First, although DistR permits the use of unsafe code, its consistency guarantees are only applicable to safe Rust code. In unsafe code blocks, developers are responsible for ensuring consistency themselves. Second, DistR’s superior performance relies on SWMR semantics exposed by applications. In cases where data is mostly under shared states (such as Mutex), DistR degenerates into a traditional DSM system; all concurrent accesses to the same data have to be centralized and serialized by the server responsible for the shared states. However, such scenarios contradict Rust’s recommended programming practices. Finally, the current implementation of DistR does not support address space layout randomization (ASLR) yet, and we have temporarily disabled it. However, DistR’s design is compatible with ASLR as long as DistR threads share the same randomized address space layout on each server. This can be achieved by delegating the randomization of stack and heap address allocation in DistR to its global controller, a feature that will be supported in future versions of DistR.
7 Evaluation
Setup. We evaluated our system on an 8-node cluster, where each node was equipped with dual Intel Xeon E5-2640 v3 processors (16 cores), 128GB of RAM, and a 40 Gbps Mellanox ConnectX-3 InfiniBand network adapter, connected by a Mellanox 100 Gbps InfiniBand switch. All servers ran Ubuntu 18.04 with kernel 5.14. We disabled hyperthreading, CPU frequency scaling, OS security mitigations in accordance with common practices [69, 72].
Methodology. We compared DistR with two state-of-the-art DSM systems, GAM [14] and Grappa [60]. For a fair comparison, we ported the evaluated applications to each baseline system and invested extensive effort in tuning parameters to achieve their best possible performance. GAM offers ordinary object read/write interfaces, and we exported it as a library to Rust and hooked pointer dereferencing to use GAM’s API without program modification. Grappa, in contrast, offers a drastically different programming abstraction that requires rewriting the program to access shared memory via delegation. Therefore, we re-implemented applications in C++ and re-structured them using Grappa’s abstractions.
7.1 Applications
We evaluated four representative datacenter applications covering a wide range of use cases and resource demands, including data analytics, microservices, scientific computation, and key-value storage, as shown in Table 1.
Application Dataset Memory (GB) Comp. Intensity (cycles/byte) DataFrame [67] h2oai [38] 64 110.13 SocialNet [33] Socfb-Penn94 [71] 64 86.09 GEMM [11] LAPACK[2] 96 300.63 KV Store [14] YCSB [22] 48 48.15
Table 1: Applications used in the evaluation.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) DataFrame (318 s) (b) SocialNet (120 ops/s) (c) GEMM (1039 s) (d) KV Store (2.7 Mops/s)
Figure 5: Application throughput when running with DistR, GAM, and Grappa, normalized to the throughput of their original implementation running on a single node. The number in the parenthesis is the original application’s throughput on a single node.
DataFrame is an in-memory data analytics framework similar to Spark [90] and Pandas [89]. We built our library atop Polars [67], a native DataFrame engine in Rust offering OLAP query APIs such as filter, groupby, and join. DataFrame organizes the dataset as columnar format tables in shared memory, and user queries will manipulate table columns by reading/writing rows and transforming them into new tables. DataFrame exploits data-level parallelism by internally partitioning columns by row into an array of small chunks where each chunk can be processed independently. We additionally applied TBox to annotate chunks from the same table column for co-location and used spawn_to to offload columnar operations to the data side to improve data locality and performance. Note that such annotations were not necessary for the application to run; they were added for additional performance optimizations.
SocialNet is a twitter-like web service from the DeathStarBench suite [33]. It is composed of 12 microservices with complicated call dependencies. Each microservice in SocialNet can scale independently with replicas, thereby offering higher throughput with more servers. SocialNet decouples the process of user texts, media resources, and storage into different microservices, and it employs RPCs to pass values (texts, media files, etc.) between them. DistR enables SocialNet to pass only references in RPCs, eliminating the serialization/deserialization overhead and redundant data copies. Because SocialNet was implemented in C++ and deployed with Docker Swarm [29], we ported it into Rust for our evaluation. We followed its original microservice structure but changed the RPC call sites to pass references instead of values, and we followed the original orchestration configuration to spread and scale each microservice in the cluster. We did not use any affinity annotations for SocialNet.
GEMM (general purpose matrix multiplication) is a highly-optimized matrix multiplication routine from the BLAS library [11]. We ported the library using the same divide-and-conquer algorithm by recursively partitioning each matrix into small chunks for parallel processing and reducing the final results. Input and output matrices are stored in the shared memory, where each subroutine will read two input matrix chunks and write the partial results back to the output matrix. Our port strictly followed the original implementation without using additional affinity annotation.
KV Store is an in-memory key-value cache engine similar to Memcached [55]. It uses a hash table to store KV pairs in shared memory and mutexes to synchronize concurrent requests. We used YCSB benchmark [22] to generate zipf load with 90% GET and 10% SET using default skewness parameter $0.99$.
7.2 Scaling Performance
In this experiment, we investigated whether DistR can speed up applications by distributing them in a cluster and how well they can scale with the number of servers used. For each application, we first ran it as is on a single server without using DSM and measured its throughput. Then, we ran the same application on DSM (subject to modifications when running Grappa) with the same configuration but on varying numbers of servers and measured the throughput normalized to its single-node throughput (i.e., strong scaling). As GAM and Grappa cannot adaptively balance the workload across servers, we evenly distributed the application’s working set and threads among all participating nodes. Ideally, an application should scale linearly and enjoy proportionally higher throughput with more nodes. However, this is usually unachievable because of the limited parallelism of real-world applications and the coherence overhead of DSM systems, and a good result for DistR will show that applications’ throughput can get close to their ideal throughput.
Figure 5 shows the results for each application respectively. DistR outperforms both GAM and Grappa in all cases. On a single node, it is 1.04–2.10$\times$ faster than two baseline DSMs, while only adding a maximum overhead of 2.42% compared to the original program. When running with multiple nodes, DistR scales up applications significantly better than GAM and Grappa. On eight nodes, DistR achieves a throughput that is 1.33–2.64$\times$ higher than that of GAM, 2.53–29.16$\times$ higher than that of Grappa.
Compared to each program’s single-machine performance, using DSM over DistR enables each program to easily leverage the available distributed resources and achieve a throughput that is 3.34–11.73$\times$ higher than their single-machine counterparts. Next, we discuss each application to explain the scalability difference between DistR and the baseline DSMs.
DataFrame. As shown in Figure 5a, compared with its original version, DataFrame running on eight nodes with DistR achieves 5.57$\times$ higher throughput, whereas with GAM and Grappa, the throughput improvements are 2.18$\times$ and 1.69$\times$, respectively. In other words, DataFrame with DistR is 2.56$\times$ and 3.29$\times$ faster than GAM and Grappa on eight nodes, respectively.
A detailed examination reveals that the performance difference comes from the shared index table in each DataFrame operation and the shared chunks between dependent DataFrame operations. In each operation, DataFrame constructs an index hash table to track the mapping from each destination chunk in the output column to all its source chunks in the input column. This index table is shared by all index-builder threads and worker threads. During processing, index-builder threads will concurrently insert into the index table using the destination chunk ID as the key and an array of source chunk IDs as the value, and worker threads will look up the shared index table and retrieve source chunks for processing. As a result, the massive writes and reads to the shared table can incur heavy coherence overhead. Further, DataFrame passes chunks as references between dependent operations and relies on the DSM system for actual data movement. However, it only performs lightweight computation over the fetched data (i.e., low compute intensity as shown in Table 1), making the coherence overhead stand out.
DistR outperforms GAM and scales much better because of its light coherence protocol, which incurs negligible object move overhead for writes and no coherence overhead for reads. The use of affinity annotations also helps DataFrame colocate worker threads with their frequently accessed data, bringing 20% additional boost (details in §7.3). GAM, in contrast, has to invalidate each cache block upon each write and read, thereby bottlenecked by the extensive coherence traffic. Grappa performs the worst in all three DSM systems due to its always-delegation programming model, which implements every global memory read/write via a delegated function call. The cost for delegation overwhelms the actual memory access latency in this case, ruining the performance of the shared hash table. Grappa’s delegation overhead actually causes a 1.23$\times$ slowdown when scaling DataFrame from a single node to two nodes.
SocialNet. Since SocialNet is microservice-based and can be deployed distributively, we added another baseline by running the original (non-DSM) code but deploying it on varying numbers of nodes. Figure 5b demonstrates the performance of all systems. SocialNet runs consistently faster with all three DSM systems compared to the original version. DistR, GAM, and Grappa achieve a 2.18$\times$, 2.02$\times$, and 1.57$\times$ speedup on a single node and a 3.51$\times$, 1.33$\times$, and 1.39$\times$ speedup on eight nodes, respectively. In the conventional setup, SocialNet requires data—such as text and media files—to be serialized into byte streams for network transmission, and then deserialized back into usable formats at the receiving end. This serialization and deserialization process is computationally intensive, particularly for large or complex data objects. In contrast, DSM systems enable SocialNet to pass references instead of the entire data values required by remote procedure calls. This approach eliminates the need for serialization and deserialization, reduces redundant data copies, and significantly enhances performance. DistR scales much better than GAM and Grappa thanks to its lightweight coherence protocol, achieving up to 2.77$\times$ and 3.16$\times$ higher throughput than GAM and Grappa, respectively.
GEMM. GEMM differs from the previous two applications in its high compute intensity and relatively infrequent shared memory accesses. In this application, matrices are transformed and divided into smaller sub-matrices for parallel processing. Each computing thread, responsible for multiplying sub-matrices, is assigned to a server. These threads cache their respective sub-matrices in the server’s local memory and access them repeatedly to compute product results. This process is highly compute-intensive. As depicted in Figure 5c, DistR and GAM scale well for GEMM and achieve 5.93$\times$, 3.82$\times$ speedup with eight nodes. In contrast, Grappa only achieves a 2.02$\times$ speedup with eight nodes due to its inability to cache sub-matrices locally, necessitating frequent remote accesses. DistR’s superior performance over GAM, with a 1.55$\times$ higher speedup on eight nodes, is primarily due to its more efficient handling of initial cross-server data accesses required when a sub-matrix is first accessed remotely. Unlike GAM, which incurs significant runtime overhead due to the maintenance of state and location of data copies, DistR directly copies data to local memory, without any complex cross-server synchronization operations, thus enhancing overall efficiency.
KV Store. KV Store is the most DSM-unfriendly application in our evaluation because it exposes poor memory locality and low compute intensity, which amplifies the overhead of cross-server memory accesses. In addition, it uses mutexes to synchronize between workers and the structure of the program does not lend itself to ownership-based read/write ordering.
Figure 5d shows the results. KV Store experiences a slowdown on all three DSM systems when scaling from a single node to two nodes (13% for DistR, 25% for GAM, and 93% for Grappa). However, the impact is mitigated when more servers are enlisted—DistR and GAM achieve 3.34$\times$ and 2.50$\times$ higher throughput on eight nodes compared to the original KV Store implementation, respectively. Due to the limited ownership semantics exposed by mutexes, DistR does not scale as well with KV Store as with other applications. DistR is 1.33$\times$ faster than GAM on eight nodes, benefiting from its adaptive load balancing and a more efficient implementation of mutexes utilizing one-sided RDMA atomic verbs, whereas GAM depends on less efficient two-sided RDMA messages for synchronization. Grappa, in contrast, incurs the highest distribution overhead and poorest scalability, primarily because each PUT/GET operation requires remote delegation, and nodes handling popular objects become bottlenecked due to skewed load.
7.3 Drill-Down Experiments
Figure 6: Effectiveness of DistR’s affinity annotations.
Latency (cycles) Average Median P90 DistR 395 356 536 Rust 364 332 496
Table 2: DistR’s Box pointer only adds a small dereferencing cost compared to Rust’s ordinary Box.
Affinity Annotations. In this experiment, we evaluated the individual contributions of affinity annotations by enabling each of them incrementally for DataFrame on eight nodes. Figure 6 reports the results. Using TBox helps DataFrame group chunks from the same column and eliminates the runtime dereference check overhead for single-column operations (e.g., filter), bringing a 12% throughput improvement. Adding spawn_to further improves the throughput by 9% by informing DistR runtime to colocate the worker thread to its input columns, which reduces cross-server memory accesses.
Runtime Dereference Checks. We measured the latency of dereferencing DistR’s Box pointer and compared it with an ordinary Rust Box pointer. Both of them point to an 8-byte object in local memory and not in CPU’s cache, which represents the common path for pointer dereferencing. Table 2 reports the results. DistR only adds a small overhead of $\sim$30 cycles. Note that this microbenchmark is extremely memory-intensive, whereas real-world applications usually employ larger object sizes and are more compute-intensive, further mitigating the runtime check overhead. For our evaluated applications, we observed a 1.02% overhead for DataFrame and a 1.14% overhead for BLAS, when they run with DistR on a single node, respectively.
[…]
Figure 7: Comparison of cache coherence costs between DistR, GAM, and Grappa on eight nodes.
Cost of Cache Coherence. In this experiment, we ran each application again on a single node and eight nodes but fixed the total amount of CPU and memory resources. For the eight-node setting, we distributed the resources evenly to each node and measured application throughput. We expect to see a slowdown due to the cost of running the coherence protocol and cross-server memory accesses, but a good result for DistR should show that application performance remains close to its original single-node version. Figure 7 reports the results. SocialNet uses pass-by-value RPCs in its original version and is significantly slower than our DSM-based version, so it is omitted in the evaluation. DistR adds only moderate cache coherence cost with an overhead of 32% in the worst case (KV Store) and 4% in the best case (GEMM). GAM and Grappa, in contrast, incur much higher overheads ranging from 10% to 98% for different applications.
8 Related Work
Software DSM Systems. Distributed cache coherence protocols and their implementations for DSM have been extensively studied since 1980s [50, 51, 17, 18, 80, 56, 32, 16, 62, 63, 61, 37]. Among them, Munin [10] and TreadMarks [6] proposed relaxed consistency models and simpler protocols trying to alleviate the coherence overhead. Recent DSM systems leveraged today’s advanced hardware such as RDMA [60, 46, 77, 14, 81, 91] to improve efficiency.
Disaggregated and Remote Memory. Memory disaggregation and remote memory techniques are another promising approach to scaling applications out of a single machine. Their key idea is to connect a host server with large memory pools [34, 47, 41] via fast datacenter network, which can be accessed by applications via OS kernel [76, 4, 86, 69] or software runtimes [73, 84, 85, 53, 35]. However, they do not provide cache coherence.
Distributed Programming Abstractions. Researchers have studied and proposed new programming languages and abstractions. Munin[10] built a type system that defines types for local and global pointers and tracks whether the pointer is shared via type checking. X10 [20, 40] and UPC [31] introduce function offloading interfaces for distributed computing and additional type annotations to reduce the runtime overhead. Ray [88] and Nu [72] are two recent systems proposing new abstractions for distributed programming. Unlike DistR, they require effort to port applications to avoid fine-grained memory sharing.
Hardware-Accelerated DSM. Specialized datacenter network technologies and emerging hardware designs stand for another trend to accelerate DSM. Clio [36], StRoM [79], and RMC [5] reduce remote memory access latency by offloading tasks into customized hardware. Concordia [87], Kona [15], and CXL [23, 24, 26, 25, 27, 91, 49, 48] enable block-level or cache-line-level memory coherence with their hardware-implemented protocols. DistR can benefit from advances in hardware support and achieve better scalability.
9 Conclusion
This paper presents DistR, a practical DSM system based on the ownership model. It automatically turns a single-machine Rust program into its distributed version with a lightweight coherence protocol guided by language semantics. DistR significantly outperforms existing state-of-the-art DSM systems, demonstrating that a language-guided DSM can achieve strong memory consistency, transparency, and efficiency simultaneously.
Acknowledgement
We thank the anonymous reviewers for their valuable and thorough comments. We are grateful to our shepherd Daniel S. Berger for his feedback. This work is supported by CNS-1763172, CNS-2007737, CNS-2006437, CNS-2106838, CNS-2147909, CNS-2128653, CNS-2301343, CNS-2330831, CNS-2403254, CNS-1764077, CNS-1956322, CNS-2106404. This work is also supported by Alibaba Group through Alibaba Research Intern Program, and funding from Amazon and Samsung.
Appendix A Artifact Appendix
A.1 Artifact Summary
DistR is an efficient, consistent, and user-friendly DSM system featuring a lightweight coherence protocol guided by language semantics. DistR allows for seamless scaling of single-machine applications to multi-server environments without sacrificing performance. Demonstrating significant improvements over existing DSM systems, DistR combines strong memory consistency, transparency, and efficiency effectively.
A.2 Artifact Check-list
  • Hardware: Intel servers equipped with InfiniBand
  • Software Environment: Rust 1.69.0, GCC 5.5, Linux Kernel 5.4, Ubuntu 18.04, MLNX-OFED 4.9
  • Public Link to Repository: https://github.com/uclasystem/DRust
  • Code License: GNU General Public License (GPL)
A.3 Description
A.3.1 DistR’s Codebase
DistR comprises four main components:
  • $\bullet$
    An RDMA communication library written in C
  • $\bullet$
    The DistR library
  • $\bullet$
    Applications integrated with DistR
  • $\bullet$
    Necessary shell scripts and configuration files
A.3.2 Deploying DistR
The initial step in deploying DistR involves cloning the source code on all involved servers: {mdframed}[style=MyFrame,nobreak=true] git clone git@github.com:uclasystem/DRust.git
Adjust several configurations according to your server setup and operational requirements:
  1. 1.
    Set the Number of Servers:
    • Define TOTAL_NUM_SERVERS in comm-lib/rdma-common.h based on the total number of available servers.
    • Similarly, adjust NUM_SERVERS in drust/src/conf.rs.
  2. 2.
    Configure the Distributed Heap Size by setting UNIT_HEAP_SIZE_GB in drust/src/conf.rs to the required heap size per server, e.g., 16 for 16GB.
  3. 3.
    Update the InfiniBand IP addresses and ports in comm-lib/rdma-server-lib.c: {mdframed}[style=MyFrame,nobreak=true] const char *ip_str[2] = {"10.0.0.1", "10.0.0.2"};
    const char *port_str[2] = {"9400", "9401"};
  4. 4.
    In drust.json, update each server’s IP address and specify three available ports.
Following configuration, build DistR as follows: {mdframed}[style=MyFrame,nobreak=true] # Compile the communication static library
cd comm-lib
make -j lib
# Copy the static library to the DistR directory
cp libmyrdma.a ../drust/
# Compile the Rust components
cd ../drust
cargo build --release
Deploy the compiled binary across all servers post-build, ensuring its correct distribution: {mdframed}[style=MyFrame,nobreak=true] scp target/release/drust user@ip:DRust/drust.out
A.3.3 Running Applications
DistR is bundled with four example applications: Dataframe, GEMM, KVStore, and SocialNet. Follow these steps to execute them:
  1. 1.
    Launch the DistR executable on all servers, excluding the main server: {mdframed}[style=MyFrame,nobreak=true] # Start the DistR process with the specified server index and application name.
    # For example, ./../drust.out -s 7 -a gemm
    cd drust
    ./../drust.out -s server_id -a app_name
  2. 2.
    On the main server: {mdframed}[style=MyFrame,nobreak=true] # Start the main DistR process with the specified application.
    cd drust
    ./../drust.out -s 0 -a app_name
More details of DistR’s installation and deployment can be found in DistR’s code repository.
Appendix B Full Coherence Protocol
This appendix provides a full version of our coherence protocol, on which our proof of memory coherence is conducted. It includes more detailed descriptions of the algorithms, specifically Algorithms 4, 6, 7, and 8, which illustrate the dereferencing and dropping procedures for each type of reference. Additionally, Algorithm 5 describes various utility functions that support these operations.
B.1 Pointer Layout
DistR extends each Rust Box pointer and reference with an additional 64-bit field, as depicted in Figure 8. Additionally, in the global address field, we reserve 16 bits as “color” bits, which are used to track the version of the referenced value. Every time a mutable reference is dereferenced for the first time, or the owner performs a write operation subsequent to the return of all immutable references, the color bits are updated. Consequently, this mechanism guarantees that any newly created immutable references will obtain a new global address [^3] In the following text, a “colored global address” means the composition of a global address and its color.. This update is critical for ensuring that outdated cache copies in the local cache are not utilized, thereby guaranteeing the coherence of DistR. Moreover, we repurpose one bit in the extension field to indicate whether the color has been updated. This design choice reduces the frequency of color updates during continuous local writes by the owner, thereby minimizing the frequency of color overflow.
B.2 Immutable Reference
The procedure for dereferencing an immutable reference is outlined in Algorithm 4. This function accepts an immutable reference, consisting of a colored global address and a local copy address (see Figure 8), and outputs a local memory address for direct reading. The initial step involves determining the location of the object using the IsLocal function (Line 4). If this address is local, DistR clears the color embedded in the global address and returns it for direct access (Line 4).
For a remote object, DistR checks whether the local copy field is null (Line 4). If this field is null, indicating no recorded local cache copy, the algorithm proceeds to examine the cache hashmap $H$ on the server executing the access. This hashmap, keyed by the colored global address, contains entries pairing local copy addresses with a count of local immutable references pointing to this copy. The algorithm checks for the presence of the colored global address $r.g$ in $H$ (Line 4). If found, it increments the reference count in that entry and updates the local copy field in the immutable reference with the local address from the entry (Lines 4-4). If $r.g$ is not in $H$, the object is copied from the remote server, stored locally, and the local address is assigned to $r.l$ (Line 4). This newly fetched object, along with its associated reference count, is then added to $H$ (Line 4). Both the update and lookup operations on $H$ are performed atomically to avoid concurrency issues.
The procedure for dropping an immutable reference is outlined in the DropRef function (Lines 4-4) in Algorithm 4. It checks the local copy address field of the reference. If it is not null, DistR decrements the reference count for this local copy in $H$. Once the reference count hits zero, the local copy becomes eligible for eviction by DistR’s runtime system.
Figure 8: DistR’s pointer layout.
Input: A shared immutable reference $r$ containing a global address $r.g$ and a local copy address $r.l$, and a local cache hashmap $\mathit{H}$.
Output: A local memory address for reading.
1
2 Function Deref($r$, $H$):
3       if IsLocal($r.g$) then
4             return ClearColor($r.g$)
5
6      else
[…]
8                   Atomic {
9                   if $r.g\in H$ then
10                         $\langle l^{\prime},cnt\rangle\leftarrow$ GetEntry($H$, $r.g$)
11                         $r.l\leftarrow l^{\prime}$
12                         UpdateEntry($H$, $r.g$, $\langle l^{\prime}$, $cnt+1\rangle$)
13
14                  else
15                         $r.l\leftarrow$ Copy(ClearColor($r.g$))
16                         InsertEntry($H$, $r.g$, $\langle r.l,1\rangle$)
17
18                  }
19
20            return $r.l$
21
22
23
24Function DropRef($r$, $H$):
25       if $r.l\neq\text{Null}$ then
26             Atomic {
27             $\langle l^{\prime}$, $cnt\rangle\leftarrow$ GetEntry($H$, $r.g$)
28             UpdateEntry($H$, $r.g$, $\langle l^{\prime}$, $cnt-1\rangle$)
29             }
30
31
Algorithm 4 Access logic for immutable reference.
1
2Function GetColor($g$):
3       return $g\gg 48$
4
5
6Function ClearColor($g$):
7       return $g$ $((1\ll 48)-1)$
8
9
10Function AppendColor($g$, $c$):
11       return ClearColor($g$) | ($c\ll 48$)
12
13
14Function ColorUpdated($o$):
15       return $(o\gg 63)$ $1$
16
17
18Function ClearUBit($o$):
19       return $o$ $((1\ll 63)-1)$
20
21
Algorithm 5 Utility functions for pointer coloring.
B.3 Mutable Reference
Input: A mutable reference $m$ containing a global address $m.g$ and the owner address $m.o$.
Output: A local memory address to be written to.
1
2 Function DerefMut($m$):
3       if IsLocal($m.g$) then
4             if $\neg$¬ColorUpdated($m.o$) then
5                   $m.o\leftarrow m.o$ | $(1\ll 63)$
6                   $c^{\prime}\leftarrow$ GetColor($m.g$) + 1
7                   $m.g\leftarrow$ AppendColor($m.g$, $c^{\prime}$)
8
9
10      else
11             $m.o\leftarrow m.o$ | $(1\ll 63)$
12             $m.g\leftarrow$ Move(ClearColor($m.g$))
13
14      return ClearColor($m.g$)
15
16
17Function DropMutRef($m$):
18       $owner\leftarrow$ ClearUBit($m.o$)
19       Write($owner$, $m.g$)
20
Algorithm 6 Access logic for mutable references.
Algorithm 6 outlines the logic for dereferencing a mutable reference consisting of a colored global address and the owner’s address (see Figure 8). It returns a local memory address for writing. The first step in the algorithm is to determine the location of the referenced object (Line 6). If the object is in the local heap partition, DistR updates the color field in the mutable reference. This update is crucial for maintaining data integrity and coherence. A critical aspect of this mechanism is the first bit of the extension field, the U bit, which indicates whether the mutable reference’s color has been updated. The algorithm checks the U bit (Line 6); if it has not been updated, DistR increments the color field and sets the U bit to 1 (Lines 6-6) . This process prevents redundant color updates during subsequent dereferences, thereby enhancing efficiency.
For remote objects, DistR moves them to the local heap partition (Line 6). The move operation involves several steps: copying the object to a local address, updating the global address field in the mutable reference to this new address, and requesting the remote server to deallocate the original object.
Upon expiration of a mutable reference, as demonstrated in the DropMutRef function (Lines 6-6), DistR writes the mutable reference’s colored global address back to the owner’s global address field. This ensures that the owner maintains an up-to-date view of the data object. Additionally, writing back the updated color guarantees that data modifications through a mutable reference on the same machine as the owner can also be reflected in owner’s global address field. This reflection changes the key used for cache look-ups by subsequent immutable references, thereby preventing access to stale data copies cached on other servers.
B.4 Owner
Input: A Box pointer $p$ containing a global address $p.g$ and a local copy address $p.l$, and a local cache hashmap $\mathit{H}$.
Output: A local memory address for reading
1 Function Deref($p$, $H$):
2       if IsLocal($p.g$) then
3             return ClearColor($p.g$)
4
5       else
[…]
7                   Atomic {
8                   if $p.g\in H$ then
9                         $\langle l^{\prime},cnt\rangle\leftarrow$ GetEntry($H$, $p.g$)
10                         $p.l\leftarrow l^{\prime}$
11                         UpdateEntry($H$, $p.g$, $\langle l^{\prime}$, $cnt+1\rangle$)
12
13                  else
14                         $p.l\leftarrow$ Copy(ClearColor($p.g$))
15                         InsertEntry($H$, $p.g$, $\langle p.l,1\rangle$)
16
17                  }
18
19            return $p.l$
20
21
Algorithm 7 Immutable access logic for owner.
Input: A Box pointer $p$ containing a global address $p.g$ and a local copy address $p.l$, and a local cache hashmap $\mathit{H}$.
Output: A local memory address to be written to
1 Function DerefMut($p$, $\mathit{H}$):
2       if IsLocal($p.g$) then
3             if $\neg$¬ColorUpdated($p.o$) then
4                   $p.o\leftarrow p.o$ | $(1\ll 63)$
5                   $c^{\prime}\leftarrow$ GetColor($p.g$) + 1
6                   $p.g\leftarrow$ AppendColor($p.g$, $c^{\prime}$)
7
8
9      else
10             if $p.g\notin H$ then
11                   $p.o\leftarrow 1\ll 63$
12                   $p.g\leftarrow$ Move(ClearColor($p.g$))
13
14            else
15                   $\langle l$, $cnt\rangle\leftarrow$ GetEntry($H$, $p.g$)
16                   RemoveEntry($H$,$p.g$)
17                   Deallocate(ClearColor(p.g))
18                   $p.g\leftarrow l$
19                   $p.o\leftarrow 1\ll 63$
20
21
22      return ClearColor($p.g$)
23
24
Algorithm 8 Mutable access logic for owner.
In DistR, memory access through an owner can be mutable or immutable. The access type can be determined at runtime, but the presence of concurrent immutable references during immutable access is not always known. Therefore, similar to immutable references, the Box pointer for an owner employs a dual-address layout. Here we discuss the handling of both immutable (read) and mutable (write) access through the owner.
  • Immutable Access. As shown in Algorithm 7, this process begins with a check of the object location. If it is in local heap partition, the global address is directly returned for dereferencing (Lines 7-7). If not, DistR emulates the handling for immutable references (Lines 7-7). The algorithm checks if the local copy address field is null. If it is, DistR checks the hashmap for existing local copies, potentially copying data to local memory, updating the hashmap, and recording the local address in the local copy field. Otherwise, DistR directly dereferences this local address.
  • Mutable Access. As detailed in Algorithm 8, writing objects through the owner begins by checking the object’s location. If it is in local heap partition and the U bit is 0 (Line 8), the owner’s color will be updated and the U bit will be set (Lines 88). If the object is remote, instead of directly moving the object to local memory, this algorithm first checks the hashmap $H$ for a local copy (Line 8). If no local copy exists, DistR mimics the mutable reference dereferencing, moving the object to a local address and updating the owner’s global address field (Line 8). If a local copy is found, DistR retrieves this local address from $H$ and sets the global address field of the mutable reference to this local address, while also deallocating the original remote copy (Lines 88).
Furthermore, when an owner goes out of scope, DistR mimics the behavior of Rust’s standard dropping function for a Box pointer. It first drops every object owned by the referenced object, and then deallocates the object pointed to by this Box pointer. Importantly, the deallocation process in DistR during moving or dropping an object also asynchronously invalidates all cached copies of the object across all servers, ensuring that references to newly allocated objects at the same address do not retrieve stale cached data.
Additionally, it is crucial to reset the $U$ bit in the extension field of the owner or a mutable reference when creating immutable references from them. This ensures that any subsequent data mutation—occurring after all immutable references have returned—can update the color again. Without this reset, subsequent immutable references following a data mutation could potentially access stale local copies.
Appendix C Memory Coherence Proof
We prove the memory coherence of DistR by demonstrating its adherence to two critical invariants as outlined in [57].
  • Single-Writer-Multiple-Reader (SWMR) Invariant. This invariant stipulates that for any specific memory location, there is either a singular writer permitted to write to it or several readers that may read from it at any given time.
  • Data-Value Invariant. This invariant asserts that if a memory location is read, it must contain the most recent value updated by the writers.
Given that the SWMR invariant is inherently upheld by Rust’s ownership-based type system, our focus shifts to proving that our algorithm guarantees the Data-Value Invariant. To accomplish this, we will demonstrate the following four invariants of our coherence protocol. Collectively, these invariants ensure the Data-Value invariant, thereby substantiating that our system achieves memory coherence in a distributed environment.
C.1 Atomicity of Global-Address Change
  • {mdframed}
    Atomicity Invariant: The global address, including the color component, stored in the Box pointer and mutable references can only be modified by a single entity at any given time.
In DistR, the colored global address is subject to modification in three specific scenarios:
  1. 1.
    Dereferencing a mutable reference. This action modifies either the global address itself or the color in the global address field of a mutable reference (Line 66 and 6 in Algorithm 6).
  2. 2.
    Mutable dereferencing of the owner. This operation either alters the color or updates the global address to point to a local address (Lines $8$ in Algorithm 8).
  3. 3.
    Dropping a mutable reference. This scenario involves updating the owner’s global address field with the colored global address stored in the mutable reference (Line 6 in Algorithm 6).
The first two scenarios impact different types of pointers (one changes the global address field of mutable references, and the other changes that of Box pointers), ensuring they do not concurrently mutate the same memory location. The same case for scenario 1 and 3. Regarding the interaction between the scenario 2 and 3, a potential overlap could occur when dropping a mutable reference created from a owner Box pointer. However, the ownership model’s single-writer property ensures that the access through the owner is prohibited while a mutable reference is alive, preventing these scenarios from happening simultaneously. Therefore, we can conclude that modifications to each colored global address are guaranteed to be atomic.
C.2 Global-Address-Change-on-Write Invariant
  • {mdframed}
    Global-Address-Change-on-Write Invariant: The colored global address of an object always changes following modifications made by either the owner or a mutable reference.
Modifications to data not present in the local heap partition inherently result in a change of the global address, as the data must be moved between servers. For data already in the local heap partition, as delineated in Lines 66 of Algorithm 6 and Lines 88 in Algorithm 8, the color in the global address field is updated. Moreover, when a mutable reference is dropped, the updated colored global address is written back to the owner (Line 6 in Algorithm 6), ensuring that the owner possesses an updated colored global address after the mutable reference alters the data.
The complexity in this invariant arises from the inclusion of the U bit, which is employed as an optimization to reduce the frequency of color updates. As detailed at the end of § B.4, every time an immutable reference is created through the owner or a mutable reference, the U bit of that entity is reset to 0. Subsequently, when mutable dereferencing occurs on a mutable reference or Box pointer, the color in that entity will be changed and the U bit will be set to 1. Future mutable dereferencing through the same owner or mutable reference will not change color again until an immutable reference is borrowed from it. This optimization does not violate this Global-Address-Change-on-Write invariant, because access to data in Rust’s ownership model can be conceptually divided into epochs. Within each epoch, if immutable references exist, shared read access is permitted through multiple immutable references and the owner. Otherwise, the epoch allows only a single entity—either the owner or a mutable reference, but not both—to read or write data. Thus, modifications to the same data within the same epoch can be regarded as a singular modification. DistR ensures that at the end of each epoch, if the data has been altered, the colored global address will have changed. As for data accesses in a single epoch, if any entity modifies the data, subsequent reads can only occur through that modifier, which can directly read the updated data and does not involve the global address at all. Consequently, the Global-Address-Change-on-Write invariant still holds with our optimization after considering modifications within the same epoch as a single write operation.
C.3 Updated-Global-Address-Visible Invariant
  • {mdframed}
    Updated-Global-Address-Visible Invariant: All subsequently created shared immutable references can observe the updated global address corresponding to the latest write operation on the data object.
Immutable references are derived from either an owner, a mutable reference, or another immutable reference. When originating from a mutable reference, the latter maintains exclusive access to the data, ensuring its global address field reflects the most recent version. Any modifications made through this mutable reference necessitate updating the global address field. Consequently, subsequent immutable references inherit this updated colored global address. In cases where an immutable reference originates from an owner, the ownership model ensures no concurrent mutable references exist for the same data. Thus, if the data has been modified through any previous mutable references, those references would have expired and updated the colored global address back to the owner. As a result, the newly created immutable reference acquires the latest colored global address. For immutable references derived from other immutable references, data mutation is prohibited during their existence. Hence, as long as the original immutable reference maintains the accurate colored global address, any new reference created from it will also possess the correct global address.
C.4 Stale-Value-Elimination Invariant
  • {mdframed}
    Stale-Value-Elimination Invariant: Following a mutation of a value through an owner or a mutable reference, stale values stored in each server’s read-only cache will not be accessed again.
This invariant can be proved using the Global-Address-Change-on-Write invariant and the Updated-Global-Address-Visible invariant. As detailed in §C.2, data access can be categorized into different epochs. During each epoch, a value is either accessible through multiple immutable references or a single mutable reference. The value may be cached on various servers in epochs of shared immutable access. Consider a scenario where we are at the end of the N-th epoch, and the value undergoes modification in the subsequent (N+1)-th epoch. The Global-Address-Change-on-Write invariant ensures that the colored global address of this value is updated. Consequently, when the local cache hashmap is next queried in future epochs via an immutable reference to this value, it will reflect the updated global address, as dictated by the Updated-Global-Address-Visible invariant. This new colored global address will differ from any prior global addresses associated with this value, which serve as keys in each server’s local cache hashmaps. Therefore, no matching entries will be found in the hashmap, effectively preventing the reading of stale values in DistR.
Finally, the validity of the Data-Value invariant is established through a proof by contradiction. If the Data-Value invariant were violated, it would imply that a stale value is read post-mutation. The stale values only exist in the local read-only cache of each server, and values can only be mutated by the owner or mutable references. This indicates that a stale value copy stored in a server’s read-only cache is accessed again after the value is mutated through the owner or mutable references, which contradicts the Stale-Value-Elimination invariant. Hence, the Data-Value invariant holds true for DistR.
By establishing both the SWMR and Data-Value invariants for DistR, we have proved the correctness of our algorithm as well as the memory coherence of DistR.
Appendix D Special Cases
D.1 Borrowing Stack Values and Partial Borrowing
Rust allows for the borrowing of stack values and partial borrowing of struct components. This leads to scenarios where mutable references may exist to stack values or a part of a struct. The challenge in these cases is that their global addresses cannot be altered when accessed via a mutable reference on a different server. To resolve this, DistR employs a copy and write-back method. It involves copying the stack value or a part of the heap struct to local memory. Upon the expiration of the mutable reference, the modified local copy is written back to the original memory location on the original server. Additionally, DistR atomically increments the color field of the struct owner from which a partial borrow occurs. For stack values, lacking a Box pointer owner, an eager eviction policy is adopted for immutable references. When the reference count of a stack value drops to 0, we eagerly delete the local cache copy and its corresponding entry on that server, ensuring that subsequent references to the same stack value will read from its original location.
D.2 Reference Creation and Ownership Transferring
During the creation of new shared immutable references from existing ones, only the global address field is duplicated, leaving the new reference’s local copy field null. This method maintains the integrity of the reference count in the hashmap and ensures that any local copy address is genuinely local. Following accesses through this new reference trigger updates to the local copy field and the hashmap by DistR.
In scenarios involving the transfer of ownership between servers, DistR deems it crucial to deallocate any local cache copies present on the source server. This step is essential to preserve consistency and prevent memory leaks during ownership transitions in distributed settings.